# Towards Responsible Evaluation for Text-to-Speech

### v1

@@ Abstract
Recent advances in text-to-speech (TTS) technology have enabled systems to produce human-indistinguishable speech, bringing benefits across accessibility, content creation, and human-computer interaction. However, current evaluation practices are increasingly inadequate for capturing the full range of capabilities, limitations, and societal implications. This position paper introduces the concept of Responsible Evaluation and argues that it is essential and urgent for the next phase of TTS development, structured through three progressive levels: (1) ensuring the faithful and accurate reflection of a model’s true capabilities, with more robust, discriminative, and comprehensive objective and subjective scoring methodologies; (2) enabling comparability, standardization, and transferability through standardized benchmarks, transparent reporting, and transferable evaluation metrics; and (3) assessing and mitigating ethical risks associated with forgery, misuse, privacy violations, and security vulnerabilities. Through this concept, we critically examine current evaluation practices, identify systemic shortcomings, and propose actionable recommendations. We hope this concept of Responsible Evaluation will foster more trustworthy and reliable TTS technology and guide its development toward ethically sound and societally beneficial applications.
Towards Responsible Evaluation for Text-to-Speech
[…]
@@ 1 Introduction
Text-to-speech (TTS) technology has made rapid progress, driven by advances in generative modeling Shen et al. (2018); Li et al. (2019); Kim et al. (2021); Ren et al. (2021); Jeong et al. (2021); Wang et al. (2023a) and the growing computational power and data Zen et al. (2019); Panayotov et al. (2015); Ma et al. (2024); Kang et al. (2024); He et al. (2024). Modern TTS systems Chen et al. (2024); Ju et al. (2024); Du et al. (2024b); Chen et al. (2025); Wang et al. (2024b) produce speech that is increasingly natural, expressive, and human-indistinguishable, offering broad benefits across accessibility, education, content creation, and voice-based human-computer interaction. At the same time, these capabilities are inherently dual-use. Realistic voice synthesis and voice cloning have already been exploited in misinformation campaigns, telecom fraud via audio deepfakes Wen et al. (2025). Moreover, biased training data can reinforce societal inequities Pinhanez et al. (2024), yielding uneven quality across demographic groups and reinforcing harmful stereotypes.
Current TTS evaluation practices, narrowly focused on technical performance in terms of naturalness, intelligibility, speaker similarity, and efficiency, have not kept pace with the growing complexity and societal reach of modern TTS applications, revealing a critical imbalance between technological advancement and its evaluation. We argue that TTS evaluation must move beyond technical performance to encompass dimensions of trustworthiness, responsibility, and ethical consideration. To this end, we put forward the concept of Responsible Evaluation for TTS, which calls for a comprehensive rethinking of how evaluation should evolve amid the rapid advancement of TTS technology across three progressive levels, and we argue that this concept is essential and urgent for the next phase of TTS development.
  • Level One: Fidelity and Accuracy. Argues that evaluation metrics reliably and faithfully reflect model capabilities;
  • Level Two: Comparability, Standardization, and Transferability. Argues that evaluation practices follow scientific rigor and fairness to enable reliable cross-system comparisons;
  • Level Three: Ethical and Risk Oversight. Argues that evaluation incorporates ethical and societal implications, aligning TTS development with the public interest and broader principles of responsible AI.
@@
Our contributions to the ongoing discourse on TTS evaluation can be summarized in three aspects: (1) comprehensive and critical diagnosis of current TTS evaluation practices. We systematically dissect standard evaluation methodologies across the TTS pipeline, covering data, training, inference, and evaluation, revealing fundamental flaws concerning fidelity, transparency, comparability, standardization, reproducibility, transferability, and ethical considerations, which collectively hinder genuine progress in TTS technology; (2) introduction and elaboration of the concept of Responsible Evaluation. We propose the concept of Responsible Evaluation built upon three progressive levels, which extend far beyond the current primary focus on technical performance, to address existing deficiencies in TTS evaluation, and align with broader responsible AI principles; (3) actionable recommendations for inspire future work on Responsible Evaluation for TTS. We articulate concrete calls to action for each level of Responsible Evaluation: (i) advancing more robust, discriminative, and comprehensive objective and subjective scoring methodologies to ensure genuine fidelity and accuracy; (ii) establishing standardized benchmarks, transparent reporting, and transferable evaluation metrics to foster comparability and transferability; and (iii) embedding systematic risk and ethical oversight, associated with forgery, misuse, privacy violations, and security vulnerabilities, aiming to steer TTS technology towards trustworthy, reliable, and societally beneficial outcomes.
@@ 2 Background: The Co-evolution of TTS Technologies and Evaluation Methods
Over the past two decades, speech synthesis has undergone a remarkable transformation, evolving from manually crafted statistical models to end-to-end deep learning systems, and more recently to approaches based on diffusion models and large language models (LLMs) Tan et al. (2021); Xie et al. (2025). Throughout this evolution, subjective evaluation has remained the foundation of TTS assessment. As new capabilities have emerged, such as zero-shot speaker adaptation and fine-grained prosody control, objective metrics like MCD and Word Error Rate (WER) have become increasingly important. These metrics provide faster and more reproducible assessments of spectral fidelity, intelligibility, and overall clarity, effectively complementing traditional subjective evaluations. As shown in Figure 1, we examine three main phases in the development of TTS technology: the statistical parametric synthesis era, the end-to-end deep learning era, and the era of diffusion models and foundation models. We analyze how evaluation methodologies have evolved alongside advances in model architectures and capabilities.
Figure 1: Evolution of TTS technology and TTS evaluation across three phases.
@@ 2.1 Statistical Parametric Synthesis Era (2000s)
Building on early rule-driven approaches Allen et al. (1987); Hallahan (1995), as well as unit selection concatenative synthesis methods Wouters and Macon (2001); Bulut et al. (2002); Stylianou (2001), the early 2000s saw the emergence of Statistical Parametric Speech Synthesis (SPSS) Yoshimura et al. (1999); Tokuda et al. (2000); Zen and Sak (2015); Fan et al. (2014). These systems model acoustic characteristics of speech such as spectral features, fundamental frequency ($F_{0}$), and duration using context-dependent HMM, DNN Zen et al. (2013), and RNN Zen and Sak (2015). The generated acoustic parameters are then passed to vocoders Kawahara (2006); Morise et al. (2016); Ai and Ling (2020) that reconstruct the speech waveform. SPSS provides a compact and flexible framework that allows precise control over prosodic elements, including pitch and timing. This makes it particularly suitable for low-resource environments and multilingual applications.
In parallel, the evaluation of TTS systems began with modest, informal approaches and has since evolved into standardized, multi-dimensional methodologies. Early research primarily relied on visual inspection of spectrograms and pitch contours, alongside informal listening tests, to assess synthesis quality Tokuda et al. (2000); Yoshimura et al. (1999). Subsequently, objective metrics such as mel-cepstral distortion (MCD), $F_{0}$ root mean square error (RMSE), and voiced/unvoiced classification error are introduced to quantitatively evaluate acoustic modeling performance Zen et al. (2013); Zen and Sak (2015). Meanwhile, subjective evaluation methods also evolved. Informal listening was gradually replaced by structured AB preference tests, enabling statistical comparisons between systems based on listener choices. Later, Mean Opinion Score (MOS) evaluations became the standard for capturing absolute judgments of naturalness on a defined scale. These methods are increasingly conducted via crowd-sourcing platforms such as Amazon Mechanical Turk, which allow for large-scale and diverse listener participation Fan et al. (2014).
@@ 2.2 Deep Learning End-to-End Era (2016-2021)
Speech synthesis technology enters a transformative era with the rise of fully neural, end-to-end architectures that significantly enhance speech naturalness and simplify the synthesis process. WaveNet Van Den Oord et al. (2016) generates high-quality raw audio by learning the long-range patterns in sound. Building on this, Tacotron Wang et al. (2017); Shen et al. (2018) uses attention-based sequence-to-sequence networks to turn text into mel-spectrograms, which a neural vocoder then converts into final waveforms. These models eliminate the need for hand-crafted linguistic features and complex alignment procedures, producing speech with more natural prosody and near-human quality. The introduction of Transformer-based models marks a further breakthrough Ren et al. (2019, 2021). In parallel, more diverse generative modeling approaches begin to emerge, including variational Lee et al. (2020); Kim et al. (2021), adversarial Ma et al. (2018), and flow-based models Miao et al. (2020); Kim et al. (2020), which unify acoustic modeling and waveform generation within a single probabilistic framework. These models reflect a broader trend toward integrated, data-driven, and highly expressive TTS systems capable of capturing the variability and richness of natural speech across different speakers, styles, and diverse contexts.
Meanwhile, TTS evaluation practice has gradually evolved to become more comprehensive, centered on subjective assessment, especially MOS, and increasingly supported by diverse objective metrics. MOS becomes the primary method for evaluating naturalness Arık et al. (2017); Gibiansky et al. (2017). Comparison MOS (CMOS) Kim et al. (2020); Li et al. (2019) and Similarity MOS (SMOS) Chen et al. (2021b) are introduced to measure relative quality and speaker similarity. Objective evaluation gains traction Kim et al. (2020) through metrics such as pitch and energy errors, and Character Error Rate (CER), which quantify acoustic fidelity and intelligibility. As non-autoregressive (NAR) systems, such as FastSpeech Ren et al. (2019) and Glow-TTS Kim et al. (2020), emerge, inference latency and model efficiency become standard evaluation criteria. Adaptation efficiency also becomes essential. Then evaluation of controllability and diversity enters an exploratory stage, with initial efforts focused on prosodic variation via pitch, duration, and sampling-based methods, though systematic evaluation metrics remain limited.
@@ 2.3 Diffusion Models and Foundation Models Era (2022-Present)
The landscape of TTS has been fundamentally transformed by the emergence of generative models Rombach et al. (2022); Ramesh et al. (2021) and LLMs OpenAI (2024); Borsos et al. (2023). Recent TTS systems leverage powerful sequence modeling to achieve unprecedented generalization, naturalness, and flexibility. Foundation models such as VALL-E Wang et al. (2023a) and its subsequent extensions Chen et al. (2024); Han et al. (2024); Du et al. (2025); Meng et al. (2025); Wang et al. (2025a); Yang et al. (2025c) redefine TTS as a conditional sequence modeling task over audio tokens, enabling zero-shot capabilities such as voice cloning and style transfer. In parallel, probabilistic generative methods, particularly diffusion models and flow-matching models, have advanced the field by enabling AR synthesis with high-fidelity Wang et al. (2025a); Jia et al. (2025), NAR synthesis with explicit duration controllability Eskimez et al. (2024); Chen et al. (2025); Wang et al. (2024b). Together, these developments mark a shift toward more unified, scalable, and general-purpose TTS systems.
The evaluation of modern TTS systems has increasingly adopted a dual-track framework that combines subjective and objective measures Wang et al. (2023a); Du et al. (2024a); Anastassiou et al. (2024). CMOS and SMOS are now widely used to assess perceived naturalness and speaker similarity, forming the core of human evaluation protocols. On the objective side, metrics such as word error rate, speaker embedding similarity, and model-based predictions of speech quality have become standard practice. Many recent approaches rely on pre-trained automatic speech recognition models Radford et al. (2022); Hsu et al. (2021); Gulati et al. (2020), speaker verification models Chen et al. (2022), and perceptual quality prediction models Baba et al. (2024); Wang et al. (2023b) to provide consistent and scalable assessments. This shift reflects a broader trend toward using neural models as evaluation tools. Despite these advances, current evaluation practices largely overlook the ethical and societal implications of highly realistic speech synthesis Lv et al. (2022); Yi et al. (2024). Issues such as identity spoofing, misinformation, and unauthorized voice replication are typically addressed only in brief disclaimers. There remains a critical need for standardized methodologies that integrate safety and misuse considerations into the core evaluation of speech generation systems.
@@ 3 Level One: Ensuring Fidelity and Accuracy in TTS Evaluation
The first level of Responsible Evaluation argues the necessity of evaluation metrics that faithfully reflect both the perceptual quality of synthesized speech and underlying system performance. When evaluation methodologies are flawed or unreliable, higher-level claims regarding comparability, standardization, or ethical considerations become unsubstantiated and ineffective. Modern TTS evaluations primarily consider dimensions including naturalness, intelligibility and robustness, speaker similarity, prosody, and system efficiency. These aspects are assessed through a combination of subjective and objective metrics. However, limitations persist in both the effectiveness of these metrics and the comprehensiveness of the evaluation dimensions covered. On the one hand, commonly used metrics sometimes fail to reflect the true capabilities of models, where objective metrics may fail to align with human perceptual judgments while subjective metrics suffer from methodological inconsistencies Chiang et al. (2023). On the other hand, the scope of evaluation dimensions remains incomplete, particularly for complex, real-world scenarios such as long-form generation and expressive content. We elaborate on these issues in the following subsections.

### v2

# Position: Towards Responsible Evaluation for Text-to-Speech

@@ Abstract
Recent advances in text-to-speech (TTS) technology have enabled systems to generate speech that is often indistinguishable from human speech, bringing benefits to accessibility, content creation, and human-computer interaction. However, current evaluation practices are increasingly inadequate for capturing the full range of capabilities, limitations, and societal impacts of modern TTS systems. This position paper introduces the concept of Responsible Evaluation and argues that it is essential and urgent for the next phase of TTS development, structured through three progressive levels: (1) ensuring the faithful and accurate reflection of a model’s true capabilities and limitations, with more robust, discriminative, and comprehensive objective and subjective scoring methodologies; (2) enabling comparability, standardization, and transferability through standardized benchmarks, transparent reporting, and transferable evaluation metrics; and (3) assessing and mitigating ethical risks associated with forgery, misuse, privacy violations, and security vulnerabilities. Through this concept, we critically examine current evaluation practices, identify systemic shortcomings, and propose actionable recommendations. We hope this concept will not only foster more reliable and trustworthy TTS technology but also guide its development toward ethically sound and societally beneficial applications.
Machine Learning, ICML
@@ 1 Introduction
Text-to-speech (TTS) has progressed rapidly in recent years, driven by advances in generative modeling (Shen et al., 2018; Kim et al., 2021; Ren et al., 2021; Jeong et al., 2021; Wang et al., 2023; Jia et al., 2025), the growing availability and diversity of speech corpora (Zen et al., 2019; Panayotov et al., 2015; Kang et al., 2024; He et al., 2024), and the scaling of computational resources. Modern TTS systems (Chen et al., 2024; Ju et al., 2024; Du et al., 2025; Hu et al., 2026) can generate speech that is highly natural, expressive, and often indistinguishable from human speech, enabling broad applications in accessibility, content creation, and human-computer interaction. At the same time, the leap in capability constitutes a double-edged sword, introducing a growing range of risks. High-fidelity voice cloning lowers the barrier to telecom fraud and disinformation through audio deepfakes. Synthetic speech further threatens biometric security, as zero-shot TTS (Wang et al., 2023) can bypass commercial automated speaker verification (ASV) systems (Wang et al., 2024b). Beyond security, biased training data can amplify societal inequities (Pinhanez et al., 2024), yielding uneven quality across demographic groups and perpetuating harmful stereotypes. More broadly, unauthorized synthesis of a person’s voice raises unresolved ethical and legal issues around consent, privacy, and digital identity.
Current TTS evaluation has not kept pace with the expanding complexity and societal reach of TTS technology. Existing practices still center on technical performance in terms of naturalness, intelligibility, speaker similarity, and efficiency, revealing a critical imbalance between technological advancement and its evaluation. We therefore argue that TTS evaluation must move beyond technical performance to encompass trustworthiness, responsibility, and ethical considerations. To this end, we put forward the concept of Responsible Evaluation for TTS, structured as three progressive levels that call for a comprehensive rethinking of how evaluation should evolve amid rapid technological progress. We argue that Responsible Evaluation is essential and urgent for the next phase of TTS development.
  • Level One: Fidelity and Accuracy. Evaluation metrics should faithfully reflect a model’s true capabilities and limitations.
  • Level Two: Comparability, Standardization, and Transferability. Evaluation practices should follow scientific rigor to enable reliable and fair cross-system comparisons.
  • Level Three: Ethical and Risk Oversight. Evaluation should incorporate ethical and societal implications, aligning TTS development with the public interest and broader principles of responsible AI.
@@
Our contributions to the discourse on TTS evaluation are threefold: (1) A comprehensive and critical diagnosis of current TTS evaluation practices. We systematically dissect standard evaluation methodologies across the TTS pipeline, covering data, training, inference, and evaluation, revealing fundamental shortcomings in fidelity, transparency, comparability, standardization, reproducibility, transferability, and ethical considerations, which collectively hinder genuine progress in TTS technology. (2) Introduction and elaboration of the concept of Responsible Evaluation. We propose a three-level concept of Responsible Evaluation that extends beyond the prevailing focus on technical performance to address structural deficiencies in current TTS evaluation and align with broader responsible AI principles. (3) Actionable recommendations for inspiring future work on responsible evaluation for TTS. We articulate concrete calls to action for each level of Responsible Evaluation: (i) advancing more robust, discriminative, and comprehensive objective and subjective scoring methodologies; (ii) establishing standardized benchmarks, transparent reporting, and transferable evaluation metrics; and (iii) integrating systematic ethical and risk oversight associated with forgery, misuse, privacy violations, and security vulnerabilities, to steer TTS technology toward trustworthy and societally beneficial outcomes.
Figure 1: Co-evolution of TTS technology and TTS evaluation across three phases.
@@ 2 Background: The Co-evolution of TTS Technologies and Evaluation Methods
Over the past two decades, speech synthesis has undergone a remarkable transformation (Tan et al., 2021a; Xie et al., 2025), evolving from manually crafted statistical models to end-to-end deep learning systems, and more recently to approaches based on diffusion models and large language models (LLMs). Throughout this evolution, subjective evaluation has remained the foundation of TTS assessment. As new capabilities have emerged, such as zero-shot speaker adaptation and fine-grained prosody control, objective metrics have become increasingly important, providing faster, reproducible assessments of specific aspects of synthesis quality and effectively complementing traditional subjective evaluations. As shown in Figure 1, we examine three main phases in the development of TTS technology: the statistical parametric synthesis era, the end-to-end deep learning era, and the era of diffusion models and foundation models. We analyze how evaluation methodologies have evolved alongside advances in model architectures and capabilities.
@@ 2.1 Statistical Parametric Synthesis Era (2000s)
Building on early rule-driven approaches (Allen et al., 1987; Hallahan, 1995), as well as unit selection concatenative synthesis methods (Moulines and Charpentier, 1990; Hunt and Black, 1996), the early 2000s saw the emergence of Statistical Parametric Speech Synthesis (SPSS) (Yoshimura et al., 1999; Tokuda et al., 2000). These systems model acoustic characteristics of speech such as spectral features, fundamental frequency ($F_{0}$), and duration using context-dependent HMM (Zen et al., 2009), and later DNN (Zen et al., 2013) and RNN (Zen and Sak, 2015). The generated acoustic parameters are then passed to signal-processing-based vocoders (Kawahara et al., 2001; Morise et al., 2016) that reconstruct the speech waveform. SPSS provides a compact and flexible framework that allows precise control over prosodic elements, including pitch and timing (Zen et al., 2009). This has led to its use in low-resource scenarios and multilingual applications (Zen et al., 2012). However, a major drawback of SPSS is its tendency to produce over-smoothed outputs (Toda et al., 2007), resulting in synthetic speech that sounds dull and lacks natural expressiveness.
In parallel, the evaluation of TTS systems began with modest, informal approaches and has since evolved into standardized, multi-dimensional methodologies. Early research (Tokuda et al., 2000; Yoshimura et al., 1999) primarily relied on visual inspection of spectrograms and pitch contours, alongside informal listening tests, to assess synthesis quality. Subsequently, objective metrics such as mel-cepstral distortion (MCD), $F_{0}$ root mean square error (RMSE), and voiced/unvoiced classification error were widely adopted to quantitatively evaluate acoustic modeling performance (Toda et al., 2007). Meanwhile, subjective evaluation methods also evolved. Informal listening was gradually replaced by structured AB preference tests (Black and Tokuda, 2005), enabling statistical comparisons between systems based on listener choices. Later, Mean Opinion Score (MOS) evaluations became the standard for capturing absolute judgments of naturalness on a defined scale (King, 2014). These methods are increasingly conducted via crowd-sourcing platforms such as Amazon Mechanical Turk, which allow for large-scale and diverse listener participation (Ribeiro et al., 2011).
@@ 2.2 Deep Learning End-to-End Era (2016-2021)
Speech synthesis technology enters a transformative era with the rise of fully neural, end-to-end architectures that significantly enhance speech naturalness and simplify the synthesis process. WaveNet (Van Den Oord et al., 2016) generates high-quality raw audio by learning the long-range patterns in sound. Building on this, Tacotron (Wang et al., 2017; Shen et al., 2018) uses attention-based sequence-to-sequence networks to turn text into mel-spectrograms, which are then transformed into waveforms by a neural vocoder. These models eliminate the need for hand-crafted linguistic features and complex alignment procedures, producing speech with more natural prosody and near-human quality. The introduction of Transformer-based models marks a further breakthrough (Li et al., 2019; Ren et al., 2019, 2021). In parallel, diverse generative modeling approaches emerged, including variational (Kim et al., 2021), adversarial (Binkowski et al., 2020), and flow-based models (Miao et al., 2020; Kim et al., 2020), culminating in models like VITS (Kim et al., 2021) that unify acoustic modeling and waveform generation within a single probabilistic framework. These innovations reflect a broader trend toward integrated, data-driven TTS systems capable of capturing the variability and richness of natural speech across diverse speakers, styles, and contexts.
Meanwhile, TTS evaluation practice has gradually evolved to become more comprehensive, centered on subjective assessment, especially MOS, and increasingly supported by diverse objective metrics. MOS becomes the primary method for evaluating naturalness (Arık et al., 2017; Gibiansky et al., 2017). Comparison MOS (CMOS) (Li et al., 2019; Kim et al., 2020) and Similarity MOS (SMOS) (Chen et al., 2021b) have become common protocols for relative naturalness and speaker similarity, respectively. Objective evaluation gains traction through metrics, including Word Error Rate (WER) for measuring robustness (Ren et al., 2019). As non-autoregressive (NAR) systems emerge, such as FastSpeech (Ren et al., 2019) and Glow-TTS (Kim et al., 2020), inference latency and model efficiency become standard evaluation criteria. Adaptation efficiency (Chen et al., 2021b) also becomes essential. Then, the evaluation of controllability and diversity enters an exploratory stage. While initial efforts on explicit prosodic modeling are quantified by pitch and energy errors (Ren et al., 2021), systematic evaluation metrics for these aspects remain limited.
@@ 2.3 Diffusion Models and Foundation Models Era (2022-Present)
The landscape of TTS has been fundamentally transformed by the emergence of generative models (Ramesh et al., 2021; Rombach et al., 2022; Borsos et al., 2023) and LLMs (OpenAI, 2024). Recent TTS systems leverage powerful sequence modeling to achieve unprecedented generalization, naturalness, and flexibility. Foundation models such as VALL-E (Wang et al., 2023) and its subsequent extensions (Chen et al., 2024; Han et al., 2024; Meng et al., 2025; Yang et al., 2025b) redefine TTS as a conditional sequence modeling task over speech tokens, enabling zero-shot capabilities such as voice cloning and style transfer. In parallel, probabilistic generative methods, particularly diffusion models (Shen et al., 2024; Ju et al., 2024) and flow-matching models (Mehta et al., 2024), have advanced the field. Breakthroughs like E2 TTS (Eskimez et al., 2024) and F5-TTS (Chen et al., 2025) demonstrate that these flow-based architectures can achieve high-fidelity, NAR synthesis with simplified alignment. Hybrid approaches such as FELLE (Wang et al., 2025a) and DiTAR (Jia et al., 2025) integrate flow matching into autoregressive frameworks in token-wise and block-wise manners, respectively, to balance long-range dependency modeling and high-fidelity speech generation. Recently, LLM-based TTS systems have gained increasing attention. Models such as CosyVoice 2/3 (Du et al., 2024b, 2025) and Qwen3-TTS (Hu et al., 2026) are initialized from LLMs (Qwen, 2025a, b) and further extended to support speech generation, leveraging the rich semantic understanding and instruction-following capabilities inherited from text modality to improve synthesis quality and enable style control within a unified foundation model framework. Together, these developments mark a shift toward unified, scalable, and general-purpose TTS systems.
The evaluation of modern TTS systems has increasingly adopted a dual-track framework that combines subjective and objective measures (Wang et al., 2023; Du et al., 2024a; Anastassiou et al., 2024). CMOS and SMOS are now widely used to assess perceived naturalness and speaker similarity, forming the core of human evaluation protocols. On the objective side, metrics such as WER, speaker embedding similarity, and model-based predictions of speech quality have become standard practice. Many recent approaches rely on pre-trained automatic speech recognition models (Radford et al., 2022; Hsu et al., 2021; Gulati et al., 2020), speaker verification models (Desplanques et al., 2020), and perceptual quality prediction models (Reddy et al., 2021; Baba et al., 2024) to provide consistent and scalable assessments. This shift reflects a broader evolution toward neural models as evaluators, culminating in the recent adoption of the LLM-as-Judges paradigm (Wang et al., 2025d, b), which goes beyond scalar scores to deliver interpretable reasoning and fine-grained quality assessments. However, current evaluation practices largely overlook the ethical and societal implications of highly realistic speech synthesis (Lv et al., 2022; Yi et al., 2024). Issues such as identity spoofing (Wang et al., 2024b), disinformation (Liu et al., 2024), and demographic bias (Pinhanez et al., 2024) are typically addressed only in ethical statements or brief disclaimers. There remains a critical need for standardized methodologies that integrate safety and misuse considerations into the core evaluation of speech generation systems.
@@ 3 Level One: Ensuring Fidelity and Accuracy in TTS Evaluation
The first level of Responsible Evaluation argues the necessity of evaluation metrics that faithfully reflect both the perceptual quality of synthesized speech and underlying system performance. When evaluation methodologies are flawed or unreliable, higher-level claims regarding comparability, standardization, or ethical considerations become unsubstantiated and ineffective. Modern TTS evaluations (Tan et al., 2021b) primarily consider dimensions including naturalness, intelligibility, robustness, speaker similarity, prosody, and system efficiency. These aspects are assessed through a combination of subjective and objective metrics. However, limitations persist in both the effectiveness of these metrics and the comprehensiveness of the evaluation dimensions covered. On the one hand, commonly used metrics sometimes fail to reflect the true capabilities of models, where objective metrics may fail to align with human perceptual judgments (Tee et al., 2026; Yang et al., 2026a) while subjective metrics suffer from methodological inconsistencies (Chiang et al., 2023). On the other hand, the scope of evaluation dimensions remains incomplete (Manku et al., 2025), particularly for complex, real-world scenarios such as long-form generation and expressive content. We elaborate on these issues in the following subsections.